# Long-term Causal Inference via Modeling Sequential Latent Confounding

### v1

@@ 6.2 Experimental Results
(a) Heterogeneous effect estimation error.
(b) Average effect estimation error.
@@
Optimal Choice of $T$ and $\mu$: We conduct experiments using different choice of $T$ and $\mu$. Specifically, to predict long-term effects, we consider the following subsets of $\mathbf{S}$: 1. $\mathbf{S}=[S_{1},S_{3}]$ with $T=2,\mu=3$; 2. $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3,\mu=2$; 3. $\mathbf{S}=[S_{1},S_{2},S_{3},S_{4},S_{5}]$ with $T=5,\mu=3$; and 4. $\mathbf{S}=[S_{1},S_{5}]$ with $T=2,\mu=1$. The results, shown in Figure 3, are sorted by the values of $\frac{\mu}{\sqrt{T-1}}$, which $3,1.41,1.34$, and $1$, respectively. As shown in Figure 3, as $\frac{\mu}{\sqrt{T-1}}$ decrease, both the estimation errors $\varepsilon_{PEHE}$ and $\varepsilon_{ATE}$ decrease. This aligns with our theoretical findings in Theorem 5.5, as a smaller $\frac{\mu}{\sqrt{T-1}}$ leads to a faster rate. Interestingly, as shown in Figure 5(b), we found that when the $\frac{\mu}{\sqrt{T-1}}$ is equal to $1$, our estimator $\hat{\tau}_{(}\mathbf{x})$ achieves lower $\varepsilon_{ATE}$ than the idealized estimator $\hat{\tau}$. This may be attributed to the fact that our estimator $\hat{\tau}(\mathbf{x})$ leverages observational data, leading to higher data efficiency.
(a) Heterogeneous effect estimation error with fixed $T$ and varying $\mu$.
(b) Average effect estimation error with fixed $T$ and varying $\mu$.
(c) Heterogeneous effect estimation error with fixed $\mu$ and varying $T$.
(d) Average effect estimation error with fixed $\mu$ and varying $T$.
@@
Vary $T$ and $\mu$: We conduct control experiments with varying $T$ and varying $\mu$ separately. First, Figures 4(a) and 4(b)) show the results for the experiments with fixed $T$ and varying $\mu$. As expected, the estimation errors on both heterogeneous and average effect estimations increase as the long-term horizon $\mu$ grows. This is primarily due to the increasing estimation errors of $\hat{f}^{\mu}(\mathbf{x})$, as established in Lemma 5.4. Second, Figures 4(c) and 4(d) show the results for the experiments with fixed $\mu$ and varying $T$. As the observed duration $T$ increases, the estimation errors decrease rapidly. It is reasonable since the larger $T$ results in a faster rate of $\hat{\tau}(\mathbf{x})$. Both findings further support our Lemma 5.4 and Theorem 5.5 again. Moreover, we observe that, with smaller $\mu$, our estimator closely matches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, showing the effectiveness of our estimator.
(a) Heterogeneous effect estimation error.
(b) Average effect estimation error.
@@
Sample sensitivity: We conduct control experiments by varying the sample sizes of both experimental and observational data. The results are presented in Figure 5. As expected, the estimator’s performance improves as the sample sizes $n_{o}$ and $n_{e}$ increase. Notably, when $n_{o}$ and $n_{e}$ become larger, our estimator $\hat{\tau}(\mathbf{x})$ closely approaches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, which again confirms the effectiveness of our estimator. Furthermore, when the sample size $n_{e}\geq 3000$ ($n_{e}\geq 6000$), the estimation errors become stable, indicating an appropriate sample size threshold.
@@ 7 Conclusion
In this paper, we extend the existing CAECB assumption (Ghassami et al., 2022) for long-term causal inference. The original assumption is restricted to settings with a one-dimensional short-term outcome. To address this limitation, we introduce a more general assumption—the functional CAECB assumption—which accommodates temporal short-term outcomes. We theoretically establish the identification result of the heterogeneous long-term causal effect under our assumption and propose a corresponding estimator by modeling sequential latent confounding. Additionally, we provide a comprehensive theoretical analysis of the estimator. Experiments confirm both the validity of our theoretical results and the effectiveness of the proposed estimator.
Impact Statement
This paper presents work whose goal is to advance the field of Long-term causal inference in the presence of latent confounders. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 3.7

### v2

@@ 6.2 Experimental Results
Comparison with Baselines We compare our method with existing methods using different choices of $T$ and $\mu$, and results are shown in Fig. 1 As expected, our method consistently outperforms all baselines, primarily because it is built upon more appropriate assumptions. In contrast, the baseline methods suffer from substantial biases resulting from incorrect assumptions. Furthermore, we observe that our estimator closely matches the idealized estimator in both magnitude and trend, highlighting the effectiveness of the proposed method.
Optimal Choice of $T$ and $\mu$ We conduct experiments using different choices of $T$ and $\mu$. Specifically, to predict long-term effects, we consider the following subsets of $\mathbf{S}$:
(a) Result of $\varepsilon_{PEHE}$.
(b) Result of $\varepsilon_{ATE}$.
@@
1. $\mathbf{S}=[S_{1},S_{3}]$ with $T=2,\mu=3$; 2. $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3,\mu=2$; 3. $\mathbf{S}=[S_{1},S_{2},S_{3},S_{4},S_{5},S_{6}]$ with $T=6,\mu=3$; and 4. $\mathbf{S}=[S_{1},S_{5}]$ with $T=2,\mu=1$. The results, shown in Figure 3, are sorted by the values of $\frac{\mu}{\sqrt{T-1}}$, which $3,1.41,1.34$, and $1$, respectively. As shown in Figure 3, as $\frac{\mu}{\sqrt{T-1}}$ decrease, both the estimation errors $\varepsilon_{PEHE}$ and $\varepsilon_{ATE}$ decrease. This aligns with our theoretical findings in Theorem 5.5, as a smaller $\frac{\mu}{\sqrt{T-1}}$ leads to a faster rate. Interestingly, as shown in Figure 3(b), we found that when the $\frac{\mu}{\sqrt{T-1}}$ is equal to $1$, our estimator $\hat{\tau}_{(}\mathbf{x})$ achieves lower $\varepsilon_{ATE}$ than the idealized estimator $\hat{\tau}$. This may be attributed to the fact that our estimator $\hat{\tau}(\mathbf{x})$ leverages observational data, leading to higher data efficiency.
(a) Fix $T$ and vary $\mu$.
(b) Fix $T$ and vary $\mu$.
(c) Fix $\mu$ and vary $T$.
(d) Fix $\mu$ and vary $T$.
@@
Vary $T$ and $\mu$ We conduct control experiments with varying $T$ and varying $\mu$ separately. First, Figures 4(a) and 4(b)) show the results for the experiments with fixed $T$ and varying $\mu$. As expected, the estimation errors on both heterogeneous and average effect estimations increase as the long-term horizon $\mu$ grows. This is primarily due to the increasing estimation errors of $\hat{f}^{\mu}(\mathbf{x})$, as established in Lemma 5.4. Second, Figures 4(c) and 4(d) show the results for the experiments with fixed $\mu$ and varying $T$. As the observed duration $T$ increases, the estimation errors decrease rapidly. It is reasonable since the larger $T$ results in a faster rate of $\hat{\tau}(\mathbf{x})$. Both findings further support our Lemma 5.4 and Theorem 5.5 again. Moreover, we observe that, with smaller $\mu$, our estimator closely matches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, showing the effectiveness of our estimator.
(a) Results of $\varepsilon_{PEHE}$.
(b) Results of $\varepsilon_{ATE}$.
@@
Sample size sensitivity We conduct control experiments by varying the sample sizes of both experimental and observational data. The results are presented in Figure 5. As expected, the estimator’s performance improves as the sample sizes $n_{o}$ and $n_{e}$ increase. Notably, when $n_{o}$ and $n_{e}$ become larger, our estimator $\hat{\tau}(\mathbf{x})$ closely approaches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, which again confirms the effectiveness of our estimator. Furthermore, when the sample size $n_{e}\geq 3000$ ($n_{o}\geq 6000$), the estimation errors become stable, indicating an appropriate sample size threshold.
@@ 7 Conclusion
In this paper, we extend the existing CAECB assumption [14] for long-term causal inference. The original assumption is restricted to settings where short- and long-term outcomes share the same scale. To address this limitation, we introduce a more general assumption, the functional CAECB assumption, which accommodates temporal short-term outcomes. Under this new assumption, we theoretically establish identification of heterogeneous long-term causal effects and practically develop an estimator by modeling sequential latent confounding. We further provide a comprehensive theoretical analysis of the estimator. Experimental results support both the theoretical findings and the estimator’s effectiveness.
Supplement to
“Long-term Causal Inference via Modeling Sequential Latent Confounding”
Appendix organization:
[…]
1
Appendix A Assumption Discussion: Latent Unconfoundedness [5], Sequential Outcomes [19], CAECB [14], FCAECB
In this section, we compare different key assumptions in [5, 19, 14] and our FCAECB assumption, and show the differences in terms of causal graphs.
Assumptions Statements Causal graphs
Latent Unconfoundedness [5] $Y(a){\perp\!\!\!\perp}A\mid\mathbf{X},\mathbf{S}(a),G=O$ [Uncaptioned image]
Sequential Outcomes [19] $(Y(a),\mathbf{S}_{3}(a)){\perp\!\!\!\perp}\mathbf{S}_{1}(a)\mid\mathbf{S}_{2}(a),\mathbf{U},\mathbf{X},G=O$ [Uncaptioned image]
CAECB [14] ${\mathbb{E}}[S(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[S(a)\mid\mathbf{X},A=1,G=O]$ $={\mathbb{E}}[Y(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[Y(a)\mid\mathbf{X},A=1,G=O]$ [Uncaptioned image]
Ours FCAECB $b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X})$ where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=0,G=O]$                                $-{\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=1,G=O]$ [Uncaptioned image]
Table 2: Various assumptions and corresponding causal graphs.
The Latent Unconfoundedness assumption states that unobserved confounders affect the long-term outcome only through short-term outcomes. In terms of causal graphs, $\mathbf{S}(a)$ fully mediates the path from $A$ to $Y(a)$, which results in ruling out the causal edge from $\mathbf{U}$ to $Y$ (Also see the discussion in [14]). This assumption ensures that $\mathbf{S}$ can serve a role of $\mathbf{U}$, making the unconfoundedness assumption hold in a certain way and thus identifying long-term effects. Instead, the Sequential Outcomes assumption and CAECB assumption allow for such a causal edge $\mathbf{U}\rightarrow Y$, while they posit different requirements. The Sequential Outcomes assumption requires that the short- and long-term outcomes can only be directly affected by the outcomes immediately preceding them, but not outcomes further in the past. In other words, $\mathbf{S}_{t+1}$ can only be affected by $\mathbf{S}_{t}$, but not by $\mathbf{S}_{t-1}$ or earlier. This assumption ensures that $\mathbf{S}_{1}$ and $\mathbf{S}_{3}$ act similarly as negative controls in proximal causal learning [29, 36], making effects identifiable. The CAECB assumption does not posit any requirement on the causal graph, but requires that the short- and long-term outcomes are measured on the same scale. Thus, the short-term outcome becomes a scale variable and may be easily violated when many short-term outcomes can be observed. This assumption states that the confounding biases in short and long term are the same, making long-term confounding bias can be identified using the observed short-term outcomes. Our FCAECB assumption likewise places no graph restriction, and further, we allow for temporal short-term outcomes instead of a single $S$. The FCAECB assumption posits that the sequence of short-term confounding biases follows a learnable temporal trend, which enables us to infer the long-term bias from observed $\mathbf{S}_{1:T}$.
Compared with the Latent Unconfoundedness and Sequential Outcomes assumptions, our proposed FCAECB assumption is neither logically stronger nor weaker. While those two assumptions impose graphical constraints, FCAECB instead posits a particular form of data-generating process. And compared to the CAECB assumption, the FCAECB assumption is weaker, as CAECB corresponds to a special case of FCAECB where $f(\mathbf{X})=1$.
Our FCAECB assumption can be further relaxed by allowing it to vary over time. In other words, we can replace $f(\mathbf{X})$ with $f(\mathbf{X},t)$ without any other modification, and the identification result also holds with almost the same proofs. However, this generalization introduces a new challenge: how to extrapolate $f(\mathbf{X},t)$, learned with $t\in\{1,...,T\}$, to $f(\mathbf{X},T+\mu)$. Addressing this challenge requires additional assumptions on $f$, such as temporal smoothness, autoregressive dependencies, etc. We leave this extension for future work.
Appendix B Broader impacts
Our work focuses on long-term causal inference, whose goal is to estimate long-term causal effects by combining long-term observational data and short-term experimental data. Our proposed method could be applied to a wide range of applications, such as decision-making in marketing and preventive measure studies in epidemiology. There are no specific possible harms or negative societal impacts that we should claim.
Appendix C Limitations
While our work relaxes some of the existing assumptions, it still relies on relatively strong ones. Regarding data requirements, similar to [5, 14, 19], our work depends on access to short-term experimental data, which may not be feasible in certain contexts, such as studies on smoking where interventions would be unethical. Additionally, the core Assumption 4.1 proposed in this paper may not hold in practice and remains untestable due to the absence of long-term outcomes in experimental data, although we propose a partial validation using short-term outcomes.
Appendix D Additional Experimental Details and Results
Data Generation Process Our data generation process partly follows [24]. Specifically, we first generate the treatments as follows: $A\mid G=O\sim\text{Bernoulli}(0.6)$ and $A\mid G=E\sim\text{Bernoulli}(0.4)$. Then we generate the observed $\mathbf{X}$ and the unobserved $\mathbf{U}$ as follows:
$(\mathbf{X},\mathbf{U})\mid A,G=E\sim\mathcal{N}([\frac{2A-1}{2},0],\begin{bmatrix}1&0\\ 0&1\end{bmatrix})$ (12)
$(\mathbf{X},\mathbf{U})\mid A,G=O\sim\mathcal{N}([\frac{1-2A}{2},0],\begin{bmatrix}1&A-0.5\\ A-0.5&1\end{bmatrix}).$
Finally, we generate the $T$-step short-term outcomes $\mathbf{S}$ and the long-term outcome $Y=S_{T+\mu}$ satisfying Assumption 4.1 as follows:
$S_{t}(A)=A+0.1A\mathbf{X}+\mathbf{X}+\mathbf{U}+\sum_{k=0}^{t}S_{k}(A)+\epsilon_{S_{t}(A)}$ (13)
where $\epsilon_{S_{t}(A)}$ are Gaussian noises. This will result in non-equal confounding bias in different time steps $t$, i.e., $\omega_{t}(\mathbf{x})=-2^{t}\mathbf{x}$.
Baselines and implementation details We compare our method with the following baselines:
  • LTEE [9]. LTEE assumes the unconfoundedness assumption, and it uses a dual-headed RNN to capture the relationships of time-dependent surrogates and then estimates the effects using the predicted long-term outcomes. We use the code available at https://github.com/GitHubLuCheng/LTEE.
  • Latent Unconfoundedness [5]. Athey et al. [5] propose a long-term effect estimator under the latent unconfoundedness assumption. We implement this method following Sec. 4.1 in [5] using correctly specified models.
  • Sequential Outcomes [19]. Imbens et al. [19] propose a long-term effect estimator by treating temporal short-term outcomes as proxies. we test three grouping strategies: (1) $S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$, (2) $S_{1},S_{\lceil T/2\rceil},S_{T}$, and (3) $S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$. We use the code available at https://github.com/CausalML/LongTermCausalInference/tree/main.
  • NNPIV [28]. Meza and Singh [28] propose an estimator for nested nonparametric instrumental variable regression, which is then applied to estimate the long-term heterogeneous effect under latent unconfoundedness and surrogacy assumption. We denote the estimators as NNPIV_LU and NNPIV_SInd under the latent unconfoundedness assumption [5] and surrogacy assumption [4], respectively. We use the code available at https://github.com/isaacmeza/NNPIV.
  • CAECB [14]. We implement CAECB [14] following the identification result in Eq. 4. Since this estimator requires only using one short-term outcome, we implement four versions, including using the first, middle, last, and random one short-term outcome among all short-term outcomes.
  • T-learner [26] trained using observational $Y$. This estimator is biased due to the latent confounding. We implement it using linear regression.
  • T-learner [26] trained using experimental $Y$. We denote this estimator as $\hat{\tau}_{exp}$. This estimator is idealized and infeasible in practice since $Y$ is missing in the experimental data and serves as an idealized benchmark. We implement it using linear regression.
  • Ours $\hat{\tau}$. We implement our method according to the algorithm described in Sec. 5.1 using correctly specified models. Unless otherwise specified, we train the model using all short-term outcomes and omit step S0.
All experiments are conducted on a machine equipped with an Intel(R) Xeon(R) E5-2620 v4 @ 2.10GHz CPU and a GeForce RTX 2080 Ti GPU. Each reported result is averaged over 50 runs.
Full Results of Table 1. We present the full result of Table 1 in Tables 4 and 3. The results are expected and consistent with the analysis presented in the main text. Our estimator outperforms all baselines and closely aligns with the idealized estimator due to the correct assumption.
Table 3: The results on synthetic data are presented as the mean and standard deviation. Lower values indicate better performance. The best-performing methods are bolded, with the exception of the idealized estimator $\hat{\tau}_{exp}$, which is excluded from the comparison.
Data $\mu=1$, $T=6$ $\mu=2$, $T=6$ $\mu=3$, $T=6$ $\mu=4$, $T=6$ $\mu=5$, $T=6$
Metrics ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$
$\hat{\tau}_{exp}$ (idealized) 3.0033 $\pm$ 1.8226 2.0722 $\pm$ 1.7681 6.0076 $\pm$ 3.6446 4.1457 $\pm$ 3.5353 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 24.0297 $\pm$ 14.5790 16.5806 $\pm$ 14.1423 48.0590 $\pm$ 29.1580 33.1620 $\pm$ 28.2845
T-learner (using obs. $Y$) 71.8431 $\pm$ 3.1612 6.8211 $\pm$ 2.8357 143.6844 $\pm$ 6.3217 13.6400 $\pm$ 5.6710 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3424 574.7384 $\pm$ 25.2877 54.5614 $\pm$ 22.6837 1149.4769 $\pm$ 50.5746 109.1205 $\pm$ 45.3685
LTEE 73.0421 $\pm$ 7.1531 10.0977 $\pm$ 6.1071 142.1183 $\pm$ 10.9446 17.3357 $\pm$ 10.0487 287.3264 $\pm$ 28.5202 32.4417 $\pm$ 19.5739 589.0638 $\pm$ 56.6554 79.3176 $\pm$ 41.2797 1168.7612 $\pm$ 122.7687 137.0583 $\pm$ 107.5567
Latent Unconfoundedness - 2.6322 $\pm$ 2.0120 - 5.2648 $\pm$ 4.0238 - 10.5293 $\pm$ 8.0471 - 21.0585 $\pm$ 16.0946 - 42.1179 $\pm$ 32.1874
Sequential Outcomes ($S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$) - 67.5881 $\pm$ 3.0107 - 135.1792 $\pm$ 6.0220 - 270.3607 $\pm$ 12.0428 - 540.7164 $\pm$ 24.0884 - 1081.4361 $\pm$ 48.1759
Sequential Outcomes ($S_{1},S_{\lceil T/2\rceil},S_{T}$) - 67.5790 $\pm$ 3.0138 - 135.1608 $\pm$ 6.0282 - 270.3237 $\pm$ 12.0553 - 540.6429 $\pm$ 24.1133 - 1081.2884 $\pm$ 48.2252
Sequential Outcomes ($S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$) - 67.5797 $\pm$ 3.0151 - 135.1622 $\pm$ 6.0309 - 270.3266 $\pm$ 12.0607 - 540.6486 $\pm$ 24.1242 - 1081.3001 $\pm$ 48.2472
NNPIV_LU 70.7124 $\pm$ 2.5320 6.1302 $\pm$ 2.8888 141.4309 $\pm$ 5.0655 12.2646 $\pm$ 5.7776 282.8595 $\pm$ 10.1320 24.5293 $\pm$ 11.5558 565.7158 $\pm$ 20.2586 49.0542 $\pm$ 23.1130 1131.4331 $\pm$ 40.5209 98.1118 $\pm$ 46.2263
NNPIV_SInd 70.7155 $\pm$ 2.5329 6.1306 $\pm$ 2.8884 141.4320 $\pm$ 5.0665 12.2626 $\pm$ 5.7782 282.8655 $\pm$ 10.1314 24.5240 $\pm$ 11.5558 565.7307 $\pm$ 20.2627 49.0492 $\pm$ 23.1115 1131.4602 $\pm$ 40.5252 98.0989 $\pm$ 46.2232
CAECB (using first $S$) 70.7257 $\pm$ 3.1153 6.7160 $\pm$ 2.7984 142.5669 $\pm$ 6.2757 13.5348 $\pm$ 5.6335 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 573.6209 $\pm$ 25.2417 54.4563 $\pm$ 22.6461 1148.3595 $\pm$ 50.5286 109.0154 $\pm$ 45.3308
CAECB (using middle $S$) 67.3754 $\pm$ 2.9827 6.3982 $\pm$ 2.6884 139.2165 $\pm$ 6.1417 13.2171 $\pm$ 5.5210 282.9014 $\pm$ 12.4626 26.8566 $\pm$ 11.1910 570.2704 $\pm$ 25.1065 54.1385 $\pm$ 22.5317 1145.0089 $\pm$ 50.3933 108.6976 $\pm$ 45.2162
CAECB (using last $S$) 36.1202 $\pm$ 2.0211 3.5778 $\pm$ 1.8603 107.9443 $\pm$ 4.9841 10.2576 $\pm$ 4.6541 251.6244 $\pm$ 11.2383 23.8972 $\pm$ 10.2153 538.9915 $\pm$ 23.8540 51.1791 $\pm$ 21.5097 1113.7291 $\pm$ 49.1287 105.7381 $\pm$ 44.1729
CAECB (using random $S$) 59.8772 $\pm$ 11.9929 5.7194 $\pm$ 2.6871 131.7174 $\pm$ 12.7309 12.5383 $\pm$ 5.3841 275.4020 $\pm$ 16.0775 26.1778 $\pm$ 10.9897 562.7700 $\pm$ 26.2888 53.4597 $\pm$ 22.2995 1137.5094 $\pm$ 50.1542 108.0188 $\pm$ 44.9689
Ours $\hat{\tau}$ 3.1110 $\pm$ 1.8230 2.1470 $\pm$ 1.7826 6.5594 $\pm$ 3.5589 4.4610 $\pm$ 3.5302 13.9945 $\pm$ 7.1682 9.2961 $\pm$ 7.0820 29.9866 $\pm$ 15.0884 19.3927 $\pm$ 14.4300 64.3076 $\pm$ 32.9868 40.5425 $\pm$ 29.7264
Table 4: The results on synthetic data are presented as the mean and standard deviation. Lower values indicate better performance. The best-performing methods are bolded, with the exception of the idealized estimator $\hat{\tau}_{exp}$, which is excluded from the comparison.
Data $\mu=5$, $T=4$ $\mu=4$, $T=5$ $\mu=3$, $T=6$ $\mu=2$, $T=7$ $\mu=1$, $T=8$
Metrics ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$
$\hat{\tau}_{exp}$ (idealized) 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709
T-learner (using obs. $Y$) 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423
LTEE 296.1849 $\pm$ 27.7377 32.9559 $\pm$ 19.0871 288.1059 $\pm$ 31.1373 34.7556 $\pm$ 20.5223 287.3264 $\pm$ 28.5202 32.4417 $\pm$ 19.5739 289.8522 $\pm$ 25.2498 41.9068 $\pm$ 25.6233 284.0581 $\pm$ 26.0439 29.2638 $\pm$ 20.0548
Latent Unconfoundedness - 10.5248 $\pm$ 8.0405 - 10.5302 $\pm$ 8.0423 - 10.5293 $\pm$ 8.0471 - 10.5270 $\pm$ 8.0476 - 10.5270 $\pm$ 8.0465
Sequential Outcomes ($S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$) - 270.3555 $\pm$ 12.0474 - 270.3607 $\pm$ 12.0428 - 270.3607 $\pm$ 12.0428 - 270.3272 $\pm$ 12.0610 - 270.3272 $\pm$ 12.0610
Sequential Outcomes ($S_{1},S_{\lceil T/2\rceil},S_{T}$) - 270.3658 $\pm$ 12.0486 - 270.3213 $\pm$ 12.0654 - 270.3237 $\pm$ 12.0553 - 270.3323 $\pm$ 12.0565 - 270.3247 $\pm$ 12.0560
Sequential Outcomes ($S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$) - 270.3554 $\pm$ 12.0473 - 270.3266 $\pm$ 12.0607 - 270.3266 $\pm$ 12.0607 - 270.3295 $\pm$ 12.0557 - 270.3295 $\pm$ 12.0557
NNPIV_LU 282.8435 $\pm$ 10.1334 24.6196 $\pm$ 11.5648 282.8424 $\pm$ 10.1261 24.5424 $\pm$ 11.5495 282.8595 $\pm$ 10.1320 24.5293 $\pm$ 11.5558 282.8690 $\pm$ 10.1315 24.5386 $\pm$ 11.5653 282.8671 $\pm$ 10.1264 24.5355 $\pm$ 11.5659
NNPIV_SInd 282.8679 $\pm$ 10.1239 24.5522 $\pm$ 11.5413 282.8408 $\pm$ 10.1337 24.5238 $\pm$ 11.5430 282.8655 $\pm$ 10.1314 24.5240 $\pm$ 11.5558 282.8683 $\pm$ 10.1288 24.5331 $\pm$ 11.5663 282.8679 $\pm$ 10.1240 24.5326 $\pm$ 11.5657
CAECB (using first $S$) 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2519 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2519 $\pm$ 12.5973 27.1744 $\pm$ 11.3048
CAECB (using middle $S$) 285.1353 $\pm$ 12.5525 27.0677 $\pm$ 11.2662 282.9014 $\pm$ 12.4625 26.8566 $\pm$ 11.1910 282.9014 $\pm$ 12.4626 26.8566 $\pm$ 11.1910 278.4314 $\pm$ 12.2812 26.4329 $\pm$ 11.0441 278.4314 $\pm$ 12.2812 26.4330 $\pm$ 11.0441
CAECB (using last $S$) 278.4314 $\pm$ 12.2812 26.4330 $\pm$ 11.0441 269.4957 $\pm$ 11.9268 25.5886 $\pm$ 10.7549 251.6244 $\pm$ 11.2383 23.8972 $\pm$ 10.2153 215.8874 $\pm$ 9.9675 20.5128 $\pm$ 9.3073 144.4755 $\pm$ 8.0850 14.3059 $\pm$ 7.4382
CAECB (using random $S$) 283.3244 $\pm$ 13.2920 26.8862 $\pm$ 11.1734 280.1459 $\pm$ 13.3821 26.5476 $\pm$ 11.0112 275.4020 $\pm$ 16.0775 26.1778 $\pm$ 10.9897 270.3812 $\pm$ 22.4630 25.5679 $\pm$ 10.7551 254.5652 $\pm$ 45.7734 24.0325 $\pm$ 10.6828
Ours $\hat{\tau}$ 19.1631 $\pm$ 12.1030 10.3532 $\pm$ 9.2908 16.6761 $\pm$ 13.1391 10.4123 $\pm$ 9.5082 13.9945 $\pm$ 7.1682 9.2961 $\pm$ 7.0820 12.7653 $\pm$ 7.4767 8.5308 $\pm$ 7.0285 12.3040 $\pm$ 7.3958 8.5348 $\pm$ 7.1581
Appendix E Proof of Theorem 3.7